OpenStudy (jazzyfa30):

please help Exponential and Logarithmic Equations solve each equation. Give an exact solution. log(base5)^(x+1)-log(base5)x=2

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

use log property : \(\large \log_a M - \log_a N = \log_a (\frac{M}{N})\)

ganeshie8 (ganeshie8):

\(\large \log_5 (x+1)-\log_5 x=2\) becomes : \(\large \log_5 \frac{x+1}{x}=2\)

OpenStudy (jazzyfa30):

\[\log_{a}b=c] b=ac right

ganeshie8 (ganeshie8):

close! but not exactly... below is the correct property : \(\large \log_a b = c \iff b = a^c\)

ganeshie8 (ganeshie8):

apply that, wat do u get ?

OpenStudy (jazzyfa30):

x+1/x=5^2

ganeshie8 (ganeshie8):

excellent ! solve \(x\)

OpenStudy (jazzyfa30):

cross multiply?????

ganeshie8 (ganeshie8):

\(\large \frac{x+1}{x} = 5^2\) \(\large \frac{x+1}{x} = 25\)

ganeshie8 (ganeshie8):

yes cross multiply

OpenStudy (jazzyfa30):

x=1/24

ganeshie8 (ganeshie8):

1/24 is \(\large \color{red}{\checkmark} \)